A room is 11 ft. by 17 ft. It needs to be painted. The cealing is 8 ft above the floor. There are two windows in the room, each one is 3 ft. by 4 ft. The door is 2.5 ft. by 6.5 ft. Find the area of the walls and ceiling.One gallon of paint covers 80.88 sq.ft. How many gallons are needed to paint the room? At $19.11 a gallon, what is the cost to paint the room?
* need to know (area of the walls and ceiling)
(number of gallons needed)
(what the cost will be)

Answer by checkley77(12844)

2*11*8+2*17*8=176+272=446 area of the walls.
11*17=187 area of the ceiling.
2(3*4)+(2.5*6.5)=2*12+16.25=24+16.25=40.25 area of the windows & the door.
446+187-40.25=592.75 ft^2 for the area to be painted.
592.75/80.88=7.329 gal. of paint will be required.
8*19.11=$152.88 cost of the paint.